# HugAgent: Evaluating LLMs in Simulating Individual-Level Human Reasoning on Open-Ended Tasks

### v2

@@ 6 Why it Happens: Ablations, Error Analysis and Mitigation Strategies
Method Belief-State Inference Belief Dynamic Update Accuracy (% $\uparrow$) Accuracy (% $\uparrow$) MAE ($\downarrow$) MSE ($\downarrow$) GPT-4o 54.44±0.06 39.30±0.03 1.93±0.10 N/A Qwen2.5-32B 60.44±0.03 38.16±0.03 2.02±0.13 N/A
Table 6: Cross-person swap test: using QA context from one participant to predict another participant’s responses in the same domain. Belief-State Inference reports Accuracy only; Belief Update reports Accuracy, MAE, and MSE. Reported as mean $\pm$ std over 5 runs and across domains.
@@ 6.2 Cross-Person Generalization
Tables 6 test how well models generalize when context is swapped across individuals.
In the cross-person setting, QA context from one participant is used to predict another participant’s stance in the same domain. Here, accuracy falls even further—e.g., GPT-4o achieves only 39.3% belief dynamics update accuracy on average across participants with an MAE of 1.93, underscoring the difficulty of transferring reasoning across individuals. This probes whether models rely on universal priors rather than individual cues. These tests show the limits of personalization: strong within-domain gains do not guarantee transfer across people, highlighting the challenge of building models that represent both individuality and general reasoning structures.
6.3 Error Analysis: Bias Patterns and Sources of Failure
Topic sensitivity.
Performance varies systematically across topics. Within-domain, belief-state inference peaks in Surveillance for both models (LLaMA 3.3 70B 80.66%; GPT-5-mini 76.89%). For update, LLaMA 3.3 70B performs best on Zoning (73.09%, MAE 1.14), whereas GPT-5-mini peaks on Healthcare (59.92%, MAE 1.37).
Cross-topic transfer sharply degrades performance. Using Surveillance context to predict Zoning, belief dynamic updates for Qwen2.5-32B-instr drop from in-domain Zoning (62.25%; MAE 1.42) to 24.84% (MAE 1.94); GPT-4o falls from 64.04% (MAE 1.26) to 42.47% (MAE 1.60).
Implication. Topic structure systematically shapes performance; adopt domain-aware prompting/retrieval/adapters and calibrate update magnitudes per domain rather than assuming uniform behavior across topics.
@@ 6.3 Error Analysis: Bias Patterns and Sources of Failure
We decompose update errors into two types: (1) wrong direction—the model predicts a belief change in the opposite direction of the human’s, and (2) change mismatch—the model fails to register that a change occurred at all. (Detailed definitions in Appendix Q.1.)
Across systems, change-mismatch errors dominate. In Healthcare, GPT-4o shows 119 mismatches vs. 5 wrong directions (DA 0.47); Qwen2.5-32B-instr records 157 vs. 5 (DA 0.31). This asymmetry holds across topics, yielding high wrong-direction accuracy ($\approx$0.93–0.98) but only moderate overall directional accuracy (0.30–0.56).
The dominance of mismatch errors indicates models preserve the sign of stance change once detected, but are insensitive to when updates should occur. Rather than inverting polarity, they often remain static despite contextual cues—suggesting weak temporal calibration or over-regularization toward “no change.” Consequently, even strong models (LLaMA 3.3 70B, GPT-4o) achieve moderate update accuracy ($\approx$60–65%) yet lag humans in Dir. Acc. by 15–20 points (Table 2). Moreover, low MAE (e.g., o3-mini, MAE 1.22) does not necessarily imply higher directional accuracy, confirming that magnitude alignment alone does not ensure correct update directionality.
Implication. LLMs are directionally stable but dynamically insensitive. Improving belief-update behavior thus requires mechanisms that better detect subtle stance shifts—context-aware memory, calibrated “no-change” thresholds, or explicit reasoning-weight tracking—rather than further tuning polarity prediction.
@@ 7 Mitigation Strategies: Insights and Guiding Principles
Our error analysis revealed structured failure modes that recur across model families. Rather than ad-hoc fixes, we position these as diagnostic insights motivating guiding principles for future work: (i) Topic sensitivity. Errors vary sharply across healthcare, zoning, and surveillance. Principle: apply calibration (offset correction, domain-specific regularization) to align predictions with human heterogeneity. (ii) Directional sensitivity. Models often preserve stance polarity but miss when updates should occur, leading to low directional accuracy. Principle: introduce temporal or contextual gating—e.g., confidence-weighted “update triggers” or explicit change-detection modules—to better capture belief shifts without overreacting to noise.
These principles underscore HugAgent’s role not only as a benchmark but also as a diagnostic tool, surfacing structured biases that standard metrics (accuracy, MAE) obscure. The larger contribution, however, is to chart a research agenda: HugAgent offers a reproducible setting in which such mitigation strategies can be systematically developed and evaluated.
@@ 8 Discussion and Open Challenges
Topic selection is not neutral. What we ask shapes who a model appears to simulate. Highly controversial topics elicit richer value trade-offs and sharper updates; homogeneous topics compress variation and inflate apparent accuracy. We propose treating topic choice as an explicit experimental variable: (i) curate pairs of domains with high vs. low opinion dispersion; (ii) report results stratified by a simple controversy index $C$ (e.g., stance variance + polarity entropy) so scores are comparable across corpora; (iii) add dilemma framings (e.g., fairness vs. safety) to probe value conflicts rather than single-axis opinions. Practically, this turns “topic selection” from a hidden confound into a controlled factor that the community can measure and replicate.
@@
Comparative benchmarks, transfer, and ethics-by-design. Personalization that succeeds only within-domain but fails across domains is not robust. We propose a comparative benchmark protocol that (i) pairs controversial vs. homogeneous topics, (ii) contrasts intuition vs. deliberation, and (iii) evaluates within-person, cross-domain transfer as a core metric. Preregistered topic panels and per-item metadata (controversy index $C$, tension flags, latency) will aid comparability. Ethically, tension induction must be consentful and minimal, with disclosed framings, capped length, and opt-out. The challenge is designing evaluation that respects persons while probing the hard cases where individuality matters most.
In short, moving from average to individual requires not only better models, but also better questions: topics that surface trade-offs, protocols that invite careful thinking, and benchmarks that reward cross-domain fidelity rather than single-domain fit.
@@ 9 Limitations and Conclusion
Our study faces constraints such as a modest human sample and potential demographic bias, though the synthetic track provides controlled variation and raw self-reports mitigate LLM cleaning effects. Even so, HugAgent offers a benchmark for average-to-individual reasoning adaptation, uniting ecological validity with scalability. Analyses expose structured failure modes in current LLMs and motivate mitigation principles, positioning HugAgent as a reproducible testbed for models that capture not only what people believe but how their reasoning evolves.
@@ Appendix A Related Work
A growing line of work simulates societies with LLM agents. Early “silicon samples” use LMs to approximate human samples in stylized tasks (Argyle et al., 2023), while Generative Agents extend to rich daily-life environments with memory and social coordination (Park et al., 2023). This has scaled to population panels—e.g., simulations of 1,000 people (Park et al., 2024) and digital-twin datasets such as Twin-2K-500 (Toubia et al., 2025)—as well as community/role-play platforms for real-time social interaction (Kaiya et al., 2023) and personification benchmarks (Xie et al., 2024b; a). These approaches provide breadth and largely static outcome measures; HugAgent adds depth via think-aloud transcripts that trace reasoning trajectories, counterfactual interventions to test belief updates, and a human test–retest reliability ceiling to anchor claims (Wurgaft et al., 2025; Nunnally & Bernstein, 1994; Cronbach, 1970).
@@
Work on Theory of Mind (ToM) in AI draws from developmental psychology tests such as the false-belief task (Wimmer & Perner, 1983), Sally-Anne (Baron-Cohen et al., 1985), and Strange Stories (Happé, 1994), later reformulated as computational tasks (Nematzadeh et al., 2018; Rabinowitz et al., 2018). Scaled language models brought ToM into broad benchmarks (Le et al., 2019; Hewitt & Cohen, 2021; Srivastava et al., 2023; Chen et al., 2024) and inspired synthetic testbeds such as BigToM (Gandhi et al., 2023), HI-TOM (He et al., 2023), FANToM (Kim et al., 2023), and MMToM-QA (Jin et al., 2024). More recent directions ground ToM in dialogues and social contexts (Chan et al., 2024; Sap et al., 2019c; Strachan et al., 2024), or frame it through Bayesian belief attribution (Ying et al., 2024). Yet these benchmarks remain synthetic, vignette-based, and decontextualized, missing ecological and demographic variability (Wang et al., 2025; Stewart et al., 2017).
Parallel lines in AI reasoning emphasize world and agent models: causal world modeling (Wong et al., 2023; Lake et al., 2017; Ellis et al., 2020) and the LAW framework, which coordinates world, agent, and language models (Hu & Shu, 2023). Within this framing, HugAgent extends ToM evaluation by asking whether models can map natural language into personalized belief states and update them consistently under interventions, bridging synthetic ToM tasks and socially grounded reasoning.
@@ Appendix C HugAgent Benchmark Details
To construct structured causal BNs from qualitative interviews, we developed a semi-structured, cognitively grounded elicitation framework. This framework guides LLMs in extracting interpretable causal structures from natural language dialogue and generating follow-up questions that balance open-ended exploration with targeted inquiry (Chickering, 2002; Pohontsch & Meyer, 2015).
@@
We asked annotators to label causal BNs using soft labels, capturing graded beliefs and allowing for variation across annotators. Our human-in-the-loop annotation tool supports annotators in assigning confidence scores to each node and edge(Zhang et al., 2023).
Annotators were also recruited from Prolific(Stewart et al., 2017). During selection, we followed three principles: (1) double-blind annotation, (2) matching annotators with similar backgrounds, and (3) using shared guidelines to maintain consistency. Annotation instructions were carefully designed to ensure reproducibility and interpretability.
@@
For the Generative Agents setting, following (Park et al., 2023), we employ Qwen2.5-32B-instruct to analyze the dialogue transcript and generate three high-level expert reflections that serve as auxiliary reasoning cues during inference. For the two retrieval-augmented variants of RAG(Lewis et al., 2020), we adopt a TF-IDF retriever to identify the top five most relevant QA pairs.
@@
We re-contacted a subset of participants for a short-interval (14-day) test–retest study. Across all sessions, 54 participants contributed data. Of these, 18 completed the retest, and 10 were retained following a demographic consistency check. We re-contacted a subset of participants for a short-interval (14-day) test–retest study. Across all sessions, 54 participants contributed data. Of these, 18 completed the retest, and 10 were retained following a demographic consistency check.
Belief State Inference yielded an accuracy of 84.49% (SD = 0.36, 95% CI: [80.82, 88.16]). By topic, accuracies were 87.50% for Surveillance (SD = 0.33, 95% CI: [81.58, 93.42]), 81.61% for Zoning (SD = 0.38, 95% CI: [74.81, 88.39]), and 84.50% for Healthcare (SD = 0.36, 95% CI: [78.25, 90.74]).
Belief Dynamic Update achieved an accuracy of 85.66% (SD = 1.27, 95% CI: [83.34, 87.70]) and a mean absolute error of 0.67 (SD = 0.83, 95% CI: [0.63, 0.73]). Across topics, accuracy was 85.75% for Healthcare (SD = 1.28), 85.33% for Surveillance (SD = 1.27), and 85.86% for Zoning (SD = 1.24). Corresponding mean absolute errors were 0.66 (SD = 0.85, 95% CI: [0.57, 0.74]) for Healthcare, 0.72 (SD = 0.74, 95% CI: [0.63, 0.80]) for Surveillance, and 0.67 (SD = 0.90, 95% CI: [0.56, 0.77]) for Zoning.
@@
Belief-State Inference Belief Dynamic Update
[…]
@@
RAG [^1] RAG (Lewis et al., 2020), RAG-FC = RAG with Full Context (Lewis et al., 2020). 71.67±0.51 78.36±1.37 76.35±0.35 50.64±0.58 53.68±0.60 51.38±0.24 1.60±0.01 1.50±0.00 1.60±0.00
RAG-FC [^1] RAG (Lewis et al., 2020), RAG-FC = RAG with Full Context (Lewis et al., 2020). 77.22±0.83 76.72±0.93 78.73±0.87 58.73±0.55 59.34±0.34 61.85±0.15 1.43±0.01 1.41±0.01 1.33±0.01
Generative Agents [^2] Generative Agents (Park et al., 2023). 76.11±0.41 77.21±0.37 75.24±0.66 57.67±0.76 55.66±0.69 61.35±0.38 1.43±0.01 1.44±0.01 1.34±0.00
Table 8: Results on Human dataset for belief-state inference and belief dynamic update tasks across three policy topics (Health, Surveillance, Zoning). Values are mean±std over 5 runs. For belief dynamic update, both Accuracy and MAE are reported separately. Best non-human results per column are in bold.
@@
Belief-State Inference Belief Dynamic Update
[…]
@@
RAG11footnotemark: 1 51.39±1.44 67.61±0.90 70.16±1.37 43.49±0.28 61.98±0.40 50.07±0.32 1.85±0.00 1.22±0.01 1.61±0.00
RAG-FC11footnotemark: 1 68.10±1.65 76.09±1.09 74.26±1.10 53.41±0.45 64.26±0.46 60.35±0.59 1.70±0.01 1.20±0.01 1.44±0.01
Generative Agents22footnotemark: 2 68.61±0.57 77.39±0.91 74.43±0.37 46.24±0.68 64.26±0.26 55.78±0.21 1.79±0.01 1.17±0.00 1.52±0.01
Table 9: Results on Synthetic dataset for belief-state inference and belief dynamic update tasks across three policy topics (Health, Surveillance, Zoning). Values are mean±std over 5 runs. For belief dynamic update, both Accuracy and MAE are reported separately. Best results per column are in bold.

### v3

@@ 5 Main Findings
Model Context Belief State Inference Belief Dynamics Update ATI (% $\uparrow$) Acc. (% $\uparrow$) Acc. (% $\uparrow$) MAE ($\downarrow$) Dir. Acc. (% $\uparrow$) GPT-4o 5 QAs 71.68±2.26 64.19±2.78 1.23±0.03 83.22±0.40 64.46±0.81 10 QAs 70.65±2.46 64.22±1.55 1.25±0.03 80.05±0.47 60.46±0.98 20+ QAs 74.66±2.86 63.11±1.19 1.29±0.04 82.27±1.02 67.29±0.83 Gemini 2.0 Flash 5 QAs 63.87±1.92 62.56±2.56 1.31±0.06 76.54±0.28 46.87±0.35 10 QAs 65.40±4.72 62.00±2.99 1.31±0.05 75.62±0.70 48.24±0.76 20+ QAs 69.95±6.01 60.55±3.19 1.35±0.03 83.31±0.19 59.76±0.21 Qwen2.5-32B-instr. 5 QAs 68.84±3.77 61.79±1.84 1.31±0.04 78.45±0.85 55.28±1.46 10 QAs 73.02±4.66 60.73±1.99 1.35±0.06 78.20±0.70 60.58±2.58 20+ QAs 77.17±0.81 58.96±2.46 1.40±0.05 76.88±0.29 64.71±0.26
Table 4: Scaling context length for both Belief State Inference and Belief Dynamics Update. Reported as mean $\pm$ std over 5 runs. Best results per column within each model group are highlighted in bold.
To examine how context length influences model performance across tasks, we varied the number of Context QAs (5, 10, 20+). Based on Table 2, we selected GPT-4o, Gemini2.0Flash, and Qwen2.5-32B-instruct as representative models from three categories: closed-source SOTA, lightweight efficient, and strong open-source baseline.
The core result is that belief state inference accuracy rises monotonically with additional dialogue and saturates at 20+ questions, whereas belief dynamics update accuracy peaks at 5–10 questions before declining. Longer context provide richer cues for recovering belief states but also introduce noise that impairs belief updating. This asymmetry suggests that context length benefits belief state inference through scale, while belief dynamics update is vulnerable to cognitive overload, likely disrupting System2 reasoning and prompting fallback to heuristic shortcutskahneman2011thinking; evans2013dual. We next investigate whether these context-length effects stem from surface-level pattern matching rather than genuine identity modeling.
6 Why It Happens: Diagnostic Ablations
Section 5 revealed a consistent failure to preserve identity across domains. Here we ask why: is it because models never learned to use personal information, or because they use it in an associative, non-generalizable way? To answer this, we conduct two diagnostic ablations.
6.1 Population Prior vs. Individual Context
Method Belief State Inference Belief Dynamics Update ATI (% $\uparrow$) Accuracy (% $\uparrow$) Accuracy (% $\uparrow$) MAE ($\downarrow$) Dir. Acc. (% $\uparrow$) GPT-4o (No-Context) 58.49±5.83 39.83±6.24 1.70±0.15 75.59±0.81 26.98±1.01 GPT-4o (Full-Context) 74.66±2.86 63.11±1.19 1.29±0.04 82.27±1.02 67.29±0.83 Qwen2.5-32B-instr. (No-Context) 50.92±6.23 32.12±3.15 1.88±0.10 69.00±1.07 6.88±0.81 Qwen2.5-32B-instr. (Full-Context) 77.17±0.81 58.96±2.46 1.40±0.05 76.88±0.29 64.71±0.26
Table 5: Comparison of No-Context (population prior) and Full-Context (with individual transcripts) settings. Reported as mean $\pm$ std over 5 runs.
We first test whether providing individual context leads to better model performance than relying solely on population-level priors. The No-Context setting provides only demographic background, while the Full-Context setting additionally includes transcripts and survey answers. As shown in Table 5, GPT-4o achieves substantially higher performance when provided with individual context: its belief-state inference accuracy increases from 58.49% with demographic priors to 74.66% with full context, and its belief-dynamics-update accuracy from 39.83% to 63.11%. The consistent gains in Full-Context indicate that models leverage individual cues rather than population priors, showing that the benchmark captures identity-sensitive reasoning rather than general demographic trends.
@@ 6.2 Cross-Person Generalization
Method Belief State Inference Belief Dynamics Update ATI (% $\uparrow$) Accuracy (% $\uparrow$) Accuracy (% $\uparrow$) MAE ($\downarrow$) Dir. Acc. (% $\uparrow$) GPT-4o (Cross-Person) 54.44±6.37 39.30±2.84 1.93±0.10 63.56±0.76 10.29±3.88 GPT-4o (Same-Person) 74.66±2.86 63.11±1.19 1.29±0.04 82.27±1.02 67.29±0.83 Qwen2.5-32B-instr. (Cross-Person) 60.44±3.22 38.16±3.28 2.02±0.13 69.17±1.06 22.15±1.24 Qwen2.5-32B-instr. (Same-Person) 77.17±0.81 58.96±2.46 1.40±0.05 76.88±0.29 64.71±0.26
Table 6: Cross-person swap test: QA context from one participant is used to predict another participant’s responses (Cross-Person) versus the same participant (Same-Person). Results are reported as mean $\pm$ std over 5 runs, averaged across domains.
We then test whether the observed gains reflect genuine identity modeling or simply the benefit of having richer, more fine-grained context. In this Cross-Person setting, QA context from one participant is used to predict another’s responses within the same domain. As shown in Table 6, performance drops sharply, e.g., GPT-4o achieves only 39.30% belief-dynamics-update accuracy with an MAE of 1.93, showing that models fail to generalize when identity cues are mismatched. These results suggest that improvements in the Full-Context setting arise from learning identity-specific patterns rather than simply benefiting from additional contextual detail.
Summary.
Together with the findings in Section 5, these ablations reveal that current models’ failure to preserve identity across domains is not because they ignore individual context in inference, but rather implies a reliance on associative context matching instead of identity-consistent reasoning.
7 Error Analysis: Bias Patterns and Sources of Failure
Domain dependence and cross-domain generalization.
Performance varies systematically across domains. Within-domain, belief state inference peaks in Surveillance for both models (LLaMA3.3-70B 80.66%; GPT-5-mini 76.89%). For belief dynamics update, LLaMA3.3-70B performs best on Zoning (73.09%, MAE 1.14), whereas GPT-5-mini peaks on Healthcare (59.92%, MAE 1.37).
Cross-domain transfer yields sharp degradation in performance. Using Surveillance context to predict Zoning, belief dynamics updates for Qwen2.5-32B-instr drop from in-domain Zoning (62.25%; MAE 1.33) to 24.84% (MAE 1.94); GPT-4o falls from 64.04% (MAE 1.26) to 42.47% (MAE 1.60).
Implication. Models do not share structural regularities in belief updating across domains, suggesting that LLMs’ belief dynamics rely more on corpus-specific semantic co-occurrence patterns than on abstract causal or cognitive mechanisms. Therefore, the degradation in cross-domain performance is not merely a domain shift issue, but rather reveals that current models lack a domain-general inductive bias for individualized human simulation.
@@ 7 Error Analysis: Bias Patterns and Sources of Failure
We decompose directional errors into two components: (1) change-detection error, reflecting the model’s failure to detect whether a belief change has occurred, and (2) direction-inference error, reflecting the model’s failure to predict the direction of that change (increase, decrease, or no change). Detailed definitions are provided in Appendix Q.1.
Across systems, change-detection errors predominate. In the Healthcare domain, GPT-4o attains a change-detection accuracy of 49.36%, compared to 88.89% for direction-inference (Dir. Acc. 77.03%). Similarly, Qwen2.5-32B-Instruct records 33.19% versus 84.29% (Dir. Acc. 68.96%). This asymmetry persists across domains, yielding consistently high accuracy in direction-inference but notably lower accuracy in change-detection, resulting in only moderate overall directional accuracy.
The prevalence of change-detection errors suggests that models often preserve the sign of stance changes once detected but fail to capture when such updates should occur. Rather than inverting polarity, they tend to remain static despite contextual cues, reflecting an over-regularization or change-averse bias toward no change. This pattern also holds for strong models such as Qwen-Max and GPT-4o, which achieve only moderate directional accuracy ($\approx$77–82%), lagging behind human performance (Dir. Acc. 88.92%) by 7–12 percentage points (Table 2). Notably, a low mean absolute error (e.g., o3-mini, MAE 1.22{1.22}) does not necessarily imply higher directional accuracy, confirming that magnitude alignment alone does not ensure correct update directionality.
Implication. LLMs tend to be directionally accurate yet change-averse, often preferring to preserve prior beliefs rather than initiating updates without strong contextual evidence. This pattern reflects an implicit bias toward stability over adaptability. It suggests that future models require mechanisms for calibrated change detection, potentially through more continuous output representations or confidence-aware update designs.
@@ 8 Mitigation Strategies: Insights and Guiding Principles
Our error analysis revealed structured failure modes that recur across model families. Rather than proposing ad-hoc fixes, we position these as diagnostic insights, each grounded in preceding analyses (Sections 67), motivating guiding principles for future work.
(i) Identity-consistent generalization gap.
Diagnosed in Section 6.2. Performance gains in Full-Context disappear under Cross-Person settings, suggesting associative matching rather than person-specific reasoning.
Principle: Learn identity-conditioned representations and persistent persona states (e.g., per-person latent slots/adapters) with cross-domain consistency regularization.
(ii) Change-aversion bias.
Evident from directional error decomposition in Section 7. Models often preserve prior beliefs even when contextual cues signal change, revealing weak mechanisms for deciding when to update beliefs.
Principle: Incorporate meta-cognitive gating, such as confidence-weighted update triggers or explicit belief-state decay, to calibrate change sensitivity and mirror human mechanisms of gradual belief revision.
(iii) Context-prioritization deficit.
Observed across context-length ablations and generalization tests (Sections 6.1). Extending context length alone does not enhance reasoning; models lack mechanisms for context prioritization and signal compression.
Principle: Future systems should implement adaptive span selection, focusing computation on high-signal evidence regions to emulate human selective attention and working-memory limits.
Synthesis.
Viewed as a whole, these principles underscore HugAgent’s role not only as a benchmark but also as a diagnostic tool, revealing structured biases that standard metrics (accuracy, MAE) often obscure. Looking ahead, HugAgent also establishes a reproducible testbed for exploring and evaluating mitigation strategies, laying the groundwork for a broader research agenda.
@@ 9 Discussion and Open Challenges
Topic selection is not neutral. Benchmarks not only measure performance, they define what counts as understanding. What we ask shapes who a model appears to simulate. Highly controversial topics elicit richer value trade-offs and sharper updates; homogeneous topics compress variation and inflate apparent accuracy. We propose treating topic choice as an explicit experimental variable: (i) curate pairs of domains with high vs. low opinion dispersion; (ii) report results stratified by a simple controversy index $C$ (e.g., stance variance + polarity entropy) so scores are comparable across corpora; (iii) add dilemma framings (e.g., fairness vs. safety) to probe value conflicts rather than single-axis opinions. Practically, this turns “topic selection” from a hidden confound into a controlled factor that the community can measure and replicate.
@@
Comparative benchmarks, transfer, and ethics-by-design. Out-of-domain generalization remains one of the most fundamental yet persistent challenges for models of human reasoning. We propose a comparative benchmark protocol that (i) pairs controversial vs. homogeneous topics, (ii) contrasts intuition vs. deliberation, and (iii) evaluates within-person, cross-domain transfer as a core metric. Preregistered topic panels and per-item metadata (controversy index $C$, tension flags, latency) will aid comparability. Ethically, tension induction must be consentful and minimal, with disclosed framings, capped length, and opt-out. The challenge is designing evaluation that respects persons while probing the hard cases where individuality matters most.
Synthesis.
Ultimately, moving reasoning from average to individual requires not only better models, but also better questions: topics that surface trade-offs, protocols that invite careful thinking, and benchmarks that reward cross-domain fidelity rather than single-domain fit.
@@ 10 Limitations and Conclusion
Our study is constrained by a modest human sample due to the resource demands of collecting deeper, higher-fidelity real human data, yet this is complemented by a synthetic track that provides controlled variation and an open, end-to-end pipeline for collecting and processing human subject data. This design enables the community to expand the dataset both in domain coverage and scale at minimal cost.
Even so, HugAgent establishes a benchmark for average-to-individual reasoning adaptation, uniting ecological validity with scalability. Across three key dimensions: (i) advancing from averaged to individualized reasoning, (ii) from behavioral mimicry to cognitive alignment, and (iii) from vignette-based to open-ended data, HugAgent redefines the methodological frontier of human behavior simulation. To our knowledge, it is the first benchmark to explicitly define and operationalize the anthropic question of how large language models simulate both human reasoning and behavior. Built as a fully open and extensible agent-simulation benchmark, HugAgent sets a new standard for evaluating how large language models simulate human-like reasoning and behavior under open-ended conditions.
@@ Appendix A Related Work
A growing line of work simulates societies with LLM agents. Early “silicon samples” use LMs to approximate human samples in stylized tasks (argyle2023out), while Generative Agents extend to rich daily-life environments with memory and social coordination (park2023thousandpeople). This has scaled to population panels—e.g., simulations of 1,000 people (park2023thousand) and digital-twin datasets such as Twin-2K-500 (toubia2025twin2k500datasetbuildingdigital)—as well as community/role-play platforms for real-time social interaction (zhao2023lyfe) and personification benchmarks (xie2024human; xie2024can). These approaches provide breadth and largely static outcome measures; HugAgent adds depth via think-aloud transcripts that trace reasoning trajectories, counterfactual interventions to test belief updates, and a human test–retest reliability ceiling to anchor claims (Wurgaft2025ScalingThinkAloud; nunnally1994psychometric; cronbach1970essentials).
@@
Work on Theory of Mind (ToM) in AI draws from developmental psychology tests such as the false-belief task (wimmer1983beliefaboutbelifs), Sally-Anne (baroncohen1985tomautism), and Strange Stories (happe1994advancedtom), later reformulated as computational tasks (nematzadeh2018evaluating; rabinowitz2018machinetheorymind). Scaled language models brought ToM into broad benchmarks (le2019revisiting; hewitt2021exploringrt; srivastava2023imitationgamequantifyingextrapolating; chen2024tombenchbenchmarkingtheorymind) and inspired synthetic testbeds such as BigToM (gandhi2023understandingsocialreasoninglanguage), HI-TOM (he2023hitombenchmarkevaluatinghigherorder), FANToM (kim2023fantombenchmarkstresstestingmachine), and MMToM-QA (jin2024mmtom). More recent directions ground ToM in dialogues and social contexts (chan2024negotiationtombenchmarkstresstestingmachine; sap2019socialiqa; strachan2024testingtom), or frame it through Bayesian belief attribution (ying2024grounding). Yet these benchmarks remain synthetic, vignette-based, and decontextualized, missing ecological and demographic variability (wang2025large; stewart2017crowdsourcing).
Parallel lines in AI reasoning emphasize world and agent models: causal world modeling (wong2022causal; lake2017building; ellis2020dreamcoder) and the LAW framework, which coordinates world, agent, and language models (hu2023languagemodelsagentmodels). Within this framing, HugAgent extends ToM evaluation by asking whether models can map natural language into personalized belief states and update them consistently under interventions, bridging synthetic ToM tasks and socially grounded reasoning.
@@ Appendix C HugAgent Benchmark Details
To construct structured causal BNs from qualitative interviews, we developed a semi-structured, cognitively grounded elicitation framework. This framework guides LLMs in extracting interpretable causal structures from natural language dialogue and generating follow-up questions that balance open-ended exploration with targeted inquiry (chickering2002optimal; pohontsch2015cognitive).
@@
We asked annotators to label causal BNs using soft labels, capturing graded beliefs and allowing for variation across annotators. Our human-in-the-loop annotation tool supports annotators in assigning confidence scores to each node and edge(zhang2023human).
Annotators were also recruited from Prolific(stewart2017crowdsourcing). During selection, we followed three principles: (1) double-blind annotation, (2) matching annotators with similar backgrounds, and (3) using shared guidelines to maintain consistency. Annotation instructions were carefully designed to ensure reproducibility and interpretability.
@@
For the Generative Agents setting, following (park2023thousandpeople), we employ Qwen2.5-32B-instruct to analyze the dialogue transcript and generate three high-level expert reflections that serve as auxiliary reasoning cues during inference. For the two retrieval-augmented variants of RAG(lewis2020retrieval), we adopt a TF-IDF retriever to identify the top five most relevant QA pairs.
@@
We re-contacted a subset of participants for a short-interval (14-day) test–retest study. Across all sessions, 54 participants contributed data. Of these, 18 completed the retest, and 13 were retained following a demographic consistency check.
Belief State Inference yielded an accuracy of 84.84% (SD = 8.90, 95% CI: [80.00, 89.68]). By topic, accuracies were 87.50% for Surveillance (SD = 8.29, 95% CI: [82.81, 92.19]), 81.54% for Zoning (SD = 15.11, 95% CI: [73.32, 89.75]), and 84.53% for Healthcare (SD = 13.90, 95% CI: [76.97, 92.09]).
Belief Dynamics Update achieved an accuracy of 85.66% (SD = 7.66, 95% CI: [80.91, 90.40]) and a mean absolute error of 0.68 (SD = 0.20, 95% CI: [0.55, 0.80]). Across topics, accuracy was 85.75% for Healthcare (SD = 10.68), 85.33% for Surveillance (SD = 12.98), and 85.86% for Zoning (SD = 10.08). Corresponding mean absolute errors were 0.66 (SD = 0.28, 95% CI: [0.49, 0.83]) for Healthcare, 0.72 (SD = 0.35, 95% CI: [0.50, 0.93]) for Surveillance, and 0.67 (SD = 0.28, 95% CI: [0.49, 0.84]) for Zoning.
In the Belief Dynamics Update tasks, we decompose directional accuracy into two components: (1) change detection—the model’s ability to detect whether a belief change has occurred, and (2) direction inference—the model’s ability to predict the direction of that change (increase, decrease, or no change) (detailed definitions are provided in Appendix Q.1).
Overall, the directional accuracy was 88.92% (SD = 9.89, 95% CI: [82.09, 96.24]). The results across topics are summarized as follows:
  • Healthcare: change detection = 80.00% (SD = 25.82, 95% CI: [61.53, 98.47]); direction inference = 96.67% (SD = 10.54, 95% CI: [89.13, 100.00]); directional accuracy = 91.67% (SD = 11.06, 95% CI: [83.76, 99.58]).
  • Surveillance: change detection = 90.00% (SD = 16.10, 95% CI: [78.48, 100.00]); direction inference = 88.33% (SD = 19.33, 95% CI: [74.51, 100.00]); directional accuracy = 88.83% (SD = 15.15, 95% CI: [77.99, 99.67]).
  • Zoning: change detection = 80.00% (SD = 23.31, 95% CI: [63.33, 96.67]); direction inference = 90.00% (SD = 21.08, 95% CI: [74.92, 100.00]); directional accuracy = 87.00% (SD = 18.14, 95% CI: [74.03, 99.97]).
@@
Belief State Inference Belief Dynamics Update
Acc. (% $\uparrow$) Acc. (% $\uparrow$) MAE ($\downarrow$)
@@
RAG [^1] RAG (lewis2020retrieval), RAG-FC = RAG with Full Context (lewis2020retrieval). 71.67±0.51 78.36±1.37 76.35±0.35 50.64±0.58 53.68±0.60 51.38±0.24 1.60±0.01 1.50±0.00 1.60±0.00
RAG-FC [^1] RAG (lewis2020retrieval), RAG-FC = RAG with Full Context (lewis2020retrieval). 77.22±0.83 76.72±0.93 78.73±0.87 58.73±0.55 59.34±0.34 61.85±0.15 1.43±0.01 1.41±0.01 1.33±0.01
Generative Agents [^2] Generative Agents (park2023thousandpeople). 76.11±0.41 77.21±0.37 75.24±0.66 57.67±0.76 55.66±0.69 61.35±0.38 1.43±0.01 1.44±0.01 1.34±0.00
[…]
@@
Belief State Inference Belief Dynamics Update
Acc. (% $\uparrow$) Acc. (% $\uparrow$) MAE ($\downarrow$)
@@
RAG [^1] RAG (lewis2020retrieval), RAG-FC = RAG with Full Context (lewis2020retrieval). 51.39±1.44 67.61±0.90 70.16±1.37 43.49±0.28 61.98±0.40 50.07±0.32 1.85±0.00 1.22±0.01 1.61±0.00
RAG-FC [^1] RAG (lewis2020retrieval), RAG-FC = RAG with Full Context (lewis2020retrieval). 68.10±1.65 76.09±1.09 74.26±1.10 53.41±0.45 64.26±0.46 60.35±0.59 1.70±0.01 1.20±0.01 1.44±0.01
Generative Agents [^2] Generative Agents (park2023thousandpeople). 68.61±0.57 77.39±0.91 74.43±0.37 46.24±0.68 64.26±0.26 55.78±0.21 1.79±0.01 1.17±0.00 1.52±0.01
[…]